OpenStudy (tay1lynn):

Simplify 5 square root of 7 end root plus 12 square root of 6 end root minus 10 square root of 7 end root minus 5 square root of 6. HELP PLEASE

OpenStudy (jocelm):

7sqrt(6)-5sqrt(7) I think is the answer to that....you really lost me.\[7\sqrt(6)-5\sqrt(7)\]

OpenStudy (jocelm):

\[5\sqrt(7)+12\sqrt(6)-10\sqrt(7)-5\sqrt(6)\]

OpenStudy (jocelm):

combine alike terms5-10=-5 and 12-5=7 so you get the answer I put in the first comment.

OpenStudy (tay1lynn):

ok im lost

OpenStudy (jocelm):

that second comment I made is the question you're asking, the 3rd comment I worked it out, in the first comment right below your question, that's the answer.

OpenStudy (jocelm):

\[7\sqrt(6)-5\sqrt(7)\]

OpenStudy (jocelm):

That is the answer....do you need help understanding how I worked it out?